an open vessel at 270C IS HEATED UNTIL3/5 OF THE AIR IN IT HAS BEEN EXPELLED.ASSUMING THE VLUME AS CONST FIND THE TEMP TO WHICH THE THE VESSEL HAS TO BE HEATED 

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Nivedhya,

Pressure=atmospheric pressure

n=PV/RT,2/5=300/T,T=750K

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Nivedhya !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR